# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ VII-B Datasets
TABLE II: Quantitative evaluations of average object motions and poses for all experiments.
@@
$00$ $01$ $02$ $03$ $04$ $05$ $06$ $18$ $20$ L1 L2 S1 S2 avg. S4U
[…]
@@
MVO $3.36$ $1.1$
DynoSAM Pose $\mathbf{1.23}$ $0.91$ $\mathbf{0.95}$ $0.27$ $0.76$ $0.56$ $2.8$ $1.15$ $0.39$ $4.2$ $0.42$ $2.88$ $1.02$ $1.6$
DynoSAM Motion $1.29$ $\mathbf{0.86}$ $1.06$ $\mathbf{0.26}$ $\mathbf{1.01}$ $0.49$ $\mathbf{0.39}$ $0.6$ $\mathbf{0.33}$ $\mathbf{0.65}$ $\mathbf{0.40}$ $\mathbf{0.52}$ $\mathbf{0.44}$ $\mathbf{0.71}$
$\text{ME}_{t}$($\mathrm{m}$) VDO-SLAM $0.11$ $0.35$ $0.43$ $\mathbf{0.15}$ $0.38$ $0.19$ $\mathbf{0.11}$ $\mathbf{0.16}$ $0.08$ $\mathbf{0.02}$
@@
MVO $0.27$ $0.03$
DynoSAM Pose $\mathbf{0.09}$ $0.40$ $0.73$ $\mathbf{0.15}$ $0.1$ $0.14$ $0.22$ $0.31$ $0.39$ $\mathbf{1.65}$ $16.6$ $0.34$ $0.64$ $0.07$
DynoSAM Motion $0.15$ $\mathbf{0.34}$ $\mathbf{0.4}$ $\mathbf{0.15}$ $\mathbf{0.09}$ $\mathbf{0.13}$ $\mathbf{0.11}$ $0.20$ $\mathbf{0.05}$ $2.19$ $\mathbf{1.86}$ $\mathbf{0.037}$ $\mathbf{0.37}$ $\mathbf{0.02}$
[…]
ClusterSLAM - - - - - - - - - - - - - $10.3$
MVO $2.7$ $\mathbf{3.13}$
DynoSAM Pose $\mathbf{1.27}$ $0.89$ $\mathbf{1.02}$ $\mathbf{0.27}$ $\mathbf{0.74}$ $\mathbf{0.6}$ $3.00$ $1.38$ $0.37$ $4.6$ $12.87$ $2.87$ $0.57$ $5.22$ $4.1$
DynoSAM Motion $1.38$ $\mathbf{0.80}$ $1.06$ $\mathbf{0.27}$ $1.04$ $0.62$ $\mathbf{2.74}$ $\mathbf{1.16}$ $\mathbf{0.33}$ $\mathbf{0.73}$ $\mathbf{0.40}$ $\mathbf{0.53}$ $\mathbf{0.26}$ $\mathbf{0.48}$ $3.2$
$\text{RPE}_{t}$($\mathrm{m}$) VDO-SLAM - - - - - - - - - - - - - - -
ClusterSLAM - - - - - - - - - - - - - $8.65$
MVO $0.28$ $0.05$
DynoSAM Pose $0.43$ $1.18$ $2.0$ $0.67$ $1.28$ $2.6$ $1.84$ $2.17$ $\mathbf{0.09}$ $\mathbf{1.05}$ $0.82$ $\mathbf{1.69}$ $\mathbf{1.26}$ $\mathbf{1.2}$ $0.06$
DynoSAM Motion $\mathbf{0.27}$ $\mathbf{0.32}$ $\mathbf{0.79}$ $\mathbf{0.19}$ $\mathbf{0.92}$ $\mathbf{0.16}$ $\mathbf{0.48}$ $\mathbf{0.2}$ $0.12$ $2.03$ $\mathbf{0.46}$ $1.92$ $1.28$ $1.42$ $\mathbf{0.04}$
To ensure valid evaluation, we pre-process these datasets offline to produce input images as well as ground truth camera and object poses/motions. For the S4U (OMD) sequence, we use color thresholding to create semantic instance masks, as the ‘square box’ class is absent from YOLO’s training data. For KITTI, we utilize pre-computed depth maps to obtain dense depth information. For all other datasets, we calculate per-pixel depth, optical flow, and semantic segmentation as outlined in Section VI-A. To maintain consistency and reproducibility, we save the processed input images to file and during evaluation load these images to guarantee identical input data for each run. To isolate the performance of our back-end estimation, we serialize the front-end output to file. This allows us to feed the same data association, measurements, and initial estimates to each DynoSAM formulation, allowing the impact of each approach to be accurately assessed. We provide custom parsers in C++ and Python to load and save the output of our front-end.
@@
TABLE III: Average percentage improvement of DynoSAM Motion estimator compared to state-of-the-art Dynamic SLAM systems. Green cell blocks indicate improvement over existing systems while red blocks indicate relatively worse accuracy.
VDO-SLAM [15] ClusterSLAM [24] MVO [17]
[…]
VII-C Camera Pose Error
The accuracy of DynoSAM’s camera pose estimation is evaluated using RPE and ATE as shown in Table I, where we outperform the state-of-the-art systems or achieve comparable results. Both DynoSAM back-ends (pose and motion estimators) demonstrate almost identical performance, which is expected as they share a common visual odometry formulation.
DynoSAM shows significant improvement over VDO-SLAM [15] on all sequences tested, and outperforms ClusterSLAM [24] in the Outdoor Cluster experiments, achieving lower average errors in both translation and rotation. On the KITTI dataset, DynoSAM performs better than DynaSLAM II [21] in relative translation and demonstrate comparable accuracy in rotation estimation. DynaSLAM II produces better results than DynoSAM in Sequences 0303030318181818 in ATE, however, the difference is minor. We believe this difference is due to the different underlying camera pose estimation algorithm employed by each system. Compared to MVO [17], we achieve similar performance on the OMD and demonstrate superior performance on KITTI.
VII-D Object Motion and Pose Errors
@@
TABLE IV: ME error per object on the OMD (S4U), comparing our motion estimator with MVO and VDO-SLAM The object ID’s reported here correspond with the following object’s in the dataset: 1 - top left, 2 - top right, 3 - bottom left, 4 - bottom right.
MVO [17] VDO-SLAM [15] DynoSAM
[…]
1 $\mathbf{0.542}$ $\mathbf{0.0169}$ $1.256$ $0.0243$ $1.138$ $0.0214$
2 $0.843$ $0.0269$ $0.770$ $0.0234$ $\mathbf{0.544}$ $\mathbf{0.0233}$
3 $1.648$ $0.0232$ $0.907$ $0.0148$ $\mathbf{0.443}$ $\mathbf{0.0086}$
4 $0.854$ $0.0309$ $0.927$ $0.0293$ $\mathbf{0.474}$ $\mathbf{0.0291}$
@@
Figure 7: Example of per-frame ME before (a) and after (b) optimization. Translation and rotation errors are shown individually; along the top and bottom rows respectively. Here we show the errors for Object $2$ (cyclist) on KITTI $00$ using the motion estimator.
[…]
Finally, Fig. 7 compares the per-frame ME before and after optimization on Object $2$ from KITTI $00$. By examining the results on a per-frame basis, we show that our framework produces accurate estimation across the entire trajectory of the object. Comparing the errors before and after optimization highlight the superior accuracy and smoothness of the back-end’s motion estimation compared to the initial-estimates from the front-end, underscoring the critical role of back-end in achieving accurate object motion estimation over the full object trajectory.
It is clear our framework is able to achieve highly accurate motion estimation in both indoor and outdoor environments and demonstrates state-of-the-art performance compared to other systems. On average, both presented back-end formulations perform comparably, however we demonstrate that the motion estimator is a more reliable method.
@@
Figure 8: Object motion errors comparing full-batch vs. sliding window. We take the $\mathbf{ME}$ for each object per frame and report the average in translation (a) and rotation (b). The motion estimator is used for both sets of results.
VII-E Sliding Window Optimisation
@@
Up to now we have presented experiment results of a full-batch solution, where all measurements are incorporated into a single optimization problem. While accurate, this approach is computationally expensive and unsuitable for online use. Furthermore, object motions often exhibit increasing independence over extended time horizons, even for objects with strong motion priors, such as vehicles on highways, thus making long-term trajectory optimization potentially unnecessary. We therefore implemented a sliding-window approach as a preliminary investigation.
@@
TABLE V: Average object ME comparison of full-batch vs. sliding window approach on the KITTI and OMD datasets.
@@
[…]
@@
This approach creates smaller, more efficient batch estimation problems every w frames, initializing overlapping variables with values from the previous window. Fig. 8 compares our full-batch and sliding window approaches, reporting the the average ME for each object per frame on the KITTI $00$ and OMD sequences. For this experiment we used a window size of $20$ frames and a $4$ frame overlap between windows, although these parameters are tunable. As shown in Table V the sliding-window method achieves comparable, albeit slightly lower, average per-frame error than full-batch on both sequences. However, full-batch demonstrates slightly more consistent accuracy, as seen in the KITTI $00$ sequence between frames $70$ and $85$, where the sliding window experiences a larger error increase at a window overlap, potentially a result of poor front-end tracking at that frame. These preliminary results suggests that incorporating motion data over a receding time horizon remains beneficial.
VII-F Computation Time
[…]
@@ VII-F Computation Time
Figure 9: Runtime breakdown of DynoSAM’s front-end. (a) shows the timing for feature tracking and (b) shows the timing per each estimation module discussed in Sections VI-B2, VI-B3 (PnP Solve), VI-B4 (Joint Optical Flow) and VI-B5 (3D Motion Refinement). Note that (b) uses log-scale for the timing axes.
VIII Downstream Tasks
@@
The proposed Dynamic SLAM system accurately estimates the motions of dynamic objects, as validated in Section VII. By eliminating the need for prior knowledge of motion models or object categories, our framework enables a range of downstream applications. This section explores how accurate motion estimation, produced by our pipeline, can facilitate dynamic object reconstruction and trajectory prediction, both of which are crucial for navigation systems.
VIII-A Dynamic Object Reconstruction

### v3

@@ VI-D Object Motion and Pose Errors
TABLE IV: Quantitative evaluations of average object motions (ME) and poses (RPE) for all experiments. Entries marked with ‘-’ indicate that results are not available for that sequence. Best results are marked in bold and second best with an underscore.
@@
$00$ $01$ $02$ $03$ $04$ $05$ $06$ $18$ 2020 L1 L2 S1 S2 avg. S4U
$\text{ME}_{r}$($\mathrm{\SIUnitSymbolDegree}$) VDO-SLAM $1.38$ $2.15$ $1.68$ $0.39$ $2.8$ $\mathbf{0.48}$ $\underline{2.8}$ $\mathbf{0.36}$ $0.47$ - - - - - $\underline{0.96}$
@@
MVO $3.36$ - - - - - - - - - - - - - $1.1$
WCPE (ours) $\mathbf{1.23}$ $\underline{0.91}$ $\mathbf{0.95}$ $\underline{0.27}$ $\mathbf{0.76}$ $0.56$ $\underline{2.8}$ $1.15$ $\underline{0.39}$ $\underline{0.86}$ $\underline{0.74}$ $\mathbf{0.68}$ $\underline{2.40}$ $\underline{1.17}$ $1.6$
WCME (ours) $\underline{1.29}$ $\mathbf{0.86}$ $\underline{1.06}$ $\mathbf{0.26}$ $\underline{1.01}$ $\underline{0.49}$ $\mathbf{0.39}$ $\underline{0.6}$ $\mathbf{0.33}$ $\mathbf{0.82}$ $\mathbf{0.70}$ $\underline{0.69}$ $\mathbf{2.36}$ $\mathbf{1.14}$ $\mathbf{0.71}$
$\text{ME}_{t}$($\mathrm{m}$) VDO-SLAM $\underline{0.11}$ $\underline{0.35}$ $\underline{0.43}$ $\mathbf{0.15}$ $0.38$ $0.19$ $\mathbf{0.11}$ $\mathbf{0.16}$ $0.57$ - - - - - $\mathbf{0.02}$
@@
MVO $0.27$ - - - - - - - - - - - - - $\underline{0.03}$
WCPE (ours) $\mathbf{0.09}$ $0.40$ $0.73$ $\mathbf{0.15}$ $\underline{0.1}$ $\underline{0.14}$ $\underline{0.22}$ $0.31$ $\underline{0.39}$ $\mathbf{0.07}$ $\underline{0.08}$ $\mathbf{0.03}$ $\mathbf{0.13}$ $\mathbf{0.08}$ $0.07$
WCME (ours) $0.15$ $\mathbf{0.34}$ $\mathbf{0.4}$ $\mathbf{0.15}$ $\mathbf{0.09}$ $\mathbf{0.13}$ $\mathbf{0.11}$ $\underline{0.20}$ $\mathbf{0.05}$ $\underline{0.08}$ $\mathbf{0.06}$ $\underline{0.04}$ $\underline{0.15}$ $\mathbf{0.08}$ $\mathbf{0.02}$
$\text{RPE}_{r}$($\mathrm{\SIUnitSymbolDegree}$) VDO-SLAM $1.40$ $1.25$ $1.34$ $0.32$ $\underline{1.04}$ $0.64$ $\mathbf{1.49}$ $\mathbf{0.38}$ $0.47$ - - - - - $3.3$
ClusterSLAM - - - - - - - - - - - - - $10.3$ -
MVO $2.7$ - - - - - - - - - - - - - $\mathbf{3.1}$
WCPE (ours) $\mathbf{1.27}$ $\underline{0.89}$ $\mathbf{1.02}$ $\mathbf{0.27}$ $\mathbf{0.74}$ $\mathbf{0.6}$ $3.00$ $1.38$ $\underline{0.37}$ $\underline{1.7}$ $\underline{2.34}$ $\mathbf{0.61}$ $\underline{2.07}$ $\underline{1.68}$ $4.1$
WCME (ours) $\underline{1.38}$ $\mathbf{0.80}$ $\underline{1.06}$ $\mathbf{0.27}$ $\underline{1.04}$ $\underline{0.62}$ $\underline{2.74}$ $\underline{1.16}$ $\mathbf{0.33}$ $\mathbf{1.3}$ $\mathbf{2.28}$ $\underline{0.67}$ $\mathbf{2.00}$ $\mathbf{1.56}$ $\underline{3.2}$
$\text{RPE}_{t}$($\mathrm{m}$) VDO-SLAM $\underline{0.28}$ $\underline{0.34}$ $\mathbf{0.30}$ $\underline{0.20}$ $\underline{0.94}$ $\underline{0.17}$ $\mathbf{0.46}$ $\mathbf{0.13}$ $\mathbf{0.09}$ - - - - - $0.06$
ClusterSLAM - - - - - - - - - - - - - $8.65$ -
MVO $\underline{0.28}$ - - - - - - - - - - - - - $\underline{0.05}$
WCPE (ours) $0.43$ $1.18$ $2.0$ $0.67$ $1.28$ $2.6$ $1.84$ $2.17$ $\mathbf{0.09}$ $\mathbf{1.84}$ $\mathbf{0.74}$ $\mathbf{0.99}$ $\underline{1.63}$ $\underline{1.33}$ $0.06$
WCME (ours) $\mathbf{0.27}$ $\mathbf{0.32}$ $\underline{0.79}$ $\mathbf{0.19}$ $\mathbf{0.92}$ $\mathbf{0.16}$ $\underline{0.48}$ $\underline{0.2}$ $\underline{0.12}$ $\underline{1.9}$ $\underline{0.72}$ $\underline{0.94}$ $\mathbf{1.50}$ $\mathbf{1.27}$ $\mathbf{0.04}$
@@
TABLE V: Average percentage improvement of DynoSAM Motion estimator compared to state-of-the-art Dynamic SLAM systems. Green cell blocks indicate improvement over existing systems while red blocks indicate relatively worse accuracy.
VDO-SLAM [zhang2020vdoslam] ClusterSLAM [Huang2019iccv] MVO [judd2024ijrr_mvo]
$\text{ME}_{r}$($\mathrm{\SIUnitSymbolDegree}$) $28\% - +$$49\%$
$\text{ME}_{t}$(m) $9\% - +$$39\%$
$\text{RPE}_{r}$($\mathrm{\SIUnitSymbolDegree}$) $9\%$ $660\%$ $9\%$
$\text{RPE}_{t}$(m) $4\% +$$681\% +$$12\%$
@@
TABLE VI: Per-object ME and ATE errors on the OMD (S4U) using the WCME. Object ID’s correspond with the following object’s in the dataset - 1: top left, 2: top right, 3: bottom left, 4: bottom right.
11 22 33 44
$\text{ME}_{r}$($\mathrm{\SIUnitSymbolDegree}$) VDO-SLAM $1.256$ $\underline{0.770}$ $\underline{0.907}$ $0.927$
MVO $\mathbf{0.542}$ $0.843$ $1.648$ $\underline{0.854}$
DynoSAM (ours) $\underline{1.138}$ $\mathbf{0.544}$ $\mathbf{0.443}$ $\mathbf{0.474}$
$\text{ME}_{t}$($\mathrm{m}$) VDO-SLAM $0.0243$ $\underline{0.0234}$ $\underline{0.0148}$ $\underline{0.0293}$
MVO $\mathbf{0.0169}$ $0.0269$ $0.0232$ $0.0309$
DynoSAM (ours) $\underline{0.0214}$ $\mathbf{0.0233}$ $\mathbf{0.0086}$ $\mathbf{0.0291}$
ATE($\mathrm{m}$) DynaSLAM II $0.41$ $0.37$ $1.09$ $0.28$
DynoSAM (ours) $\mathbf{0.09}$ $\mathbf{0.21}$ $\mathbf{0.08}$ $\mathbf{0.15}$
@@
Figure 8: Example of per-frame ME before (a) and after (b) optimization using the WCME. We show results for object 22, a non-rigid body (cyclist) on KITTI $00$. Translation and rotation errors are shown individually; along the top and bottom rows respectively. The resulting trajectories are compared to the ground truth in (c). These results demonstrate our systems capacity to handle non-rigid bodies in the case where the object experiences a predominant motion.
Table V demonstrates that DynoSAM outperforms MVO [judd2024ijrr_mvo] across all metrics, except in $\text{RPE}_{r}$ on the OMD where MVO is more accurate by a small margin. Our framework is particularly accurate in translation, with a $39\%$ improvement in MEt and $12\%$ improvement in RPEt. However, reporting the ME individually for each object in the S4U sequence in Table VI shows that that DynoSAM outperforms MVO on most objects, thereby reinforcing our accuracy even on this highly dynamic dataset. We additionally outperform VDO-SLAM on all objects. Compared to DynaSLAM II we show an average $445\%$ improvement in object ATE.
@@
Figure 9: Object motion errors comparing full-batch vs sliding window. We take the $\mathbf{ME}$ (rotation and translation) for each object per frame and report the average for KITTI $00$ (a) and OMD (S4U) (b). The WCME is used for both sets of results.
Compared to VDO-SLAM [zhang2020vdoslam], our framework in general exhibits superior accuracy in object motion and pose estimation. Such improvement is particularly pronounced on the OMD, where DynoSAM is $34\%$ more accurate in rotation and $21\%$ in translation errors, as measured by the ME metric. However, there remains some situations such as KITTI $05$, $06$ and $18$ where VDO-SLAM performs better. While the difference on KITTI $05$ is marginal, $06$ and $18$ exhibit larger performance differences that merit further investigation. Although these two systems share a similar WCME formulation and are expected to yield comparable results, they differ in the solvers they use and in their underlying approaches to object tracking. Since we report performance averaged across all objects, such variations can significantly affect the final results, especially in sequences with many objects (KITTI $18$). We believe this is a key reason for the discrepancies in the corresponding VDO-SLAM and DynoSAM performances. In the future, we will explore tracking quality by including the MOTP metric [bescos2021ral]; however, this is outside the main scope of this paper.
Finally we compare between WCME and WCPE. Across all sequences the average error difference is only $0.06\text{\,}\mathrm{m}$ in translation and $0.11\text{\,}\mathrm{\SIUnitSymbolDegree}$ in rotation, demonstrating comparable performance. Out of the $14$ sequences tested there are only three sequences (KITTI $03$, $06$ and $20$) where the difference in error is greater than one standard-deviation, calculated independently for translation and rotation across all sequences. In these instances, WCPE’s slightly degraded performance is likely due to its sensitivity to object pose initialization.
In WCPE, object poses are initialized at each time-step using the centroid of observed object points. Since these observations often represent only a partial view of the object, the resulting initialization may not be consistent with previous estimates, introducing additional noise. This limitation can be mitigated in future work by using pose prediction methods to provide better priors or serve as complementary measurements. While WCPE is not necessarily expected to outperform WCME in motion estimation specifically, it remains highly relevant, as direct pose estimation is attractive for many downstream applications such as navigation [salzmann2020trajectron++] and reconstruction [Wang2025icra_dynorecon]. Therefore, we believe that it remains relevant to introduce WCPE and examine its performance despite the slight degradation in motion accuracy.
Finally, Fig. 8 compares the per-frame ME before and after optimization on object $2$ (cyclist) from KITTI $00$, highlighting our accurate estimation across the entire trajectory. These results further demonstrate our systems capability to deal with only partially rigid objects, i.e. a cyclist, despite its motion breaking the explicit rigid-body assumption. We hypothesize that as long as the object experiences a single predominant motion, and sufficient object points can be tracked, our method will recover the overall trajectory, as shown in Fig. 8 (c).
VI-E Sliding Window Optimization
@@
We have so far presented results from a full-batch solution that incorporates all measurements into a single optimization. This solution, while accurate, is computationally intensive and unsuitable for online use. Moreover, we hypotheise that object motions should exhibit increasing independence over extended time horizons, even for objects with strong motion priors, such as vehicles on highways, thus making optimization over the global trajectory potentially unnecessary.
Therefore, we implemented a sliding-window approach that solves smaller, more efficient batch problems every $w$ frames, reusing estimates from the previous window to initialize overlapping variables. As shown in Table VII the sliding-window method achieves marginally better per-frame error. Furthermore, Fig. 9 compares the per-frame error of each approach. The average ME for each object on the KITTI $00$ and OMD sequences is reported. Note that full-batch demonstrates slightly more consistent accuracy, as seen in Fig. 9 (a) between frames $70$ and $85$. The larger error occurs at window overlap and is likely a result of poor front-end tracking at that frame. Our preliminary results indicate that incorporating motion data over a receding time horizon is beneficial.
VI-F Computation Time
@@
TABLE VII: Average object ME comparison of full-batch vs. sliding window approach on the KITTI and OMD datasets.
@@
$\text{ME}_{r}$($\mathrm{\SIUnitSymbolDegree}$) $\text{ME}_{t}$(m) $\text{ME}_{r}$($\mathrm{\SIUnitSymbolDegree}$) $\text{ME}_{t}$(m)
@@
Figure 10: Runtime breakdown of DynoSAM’s front-end. (a) shows the timing for feature tracking and (b) shows the timing per each estimation module discussed in Sections V-B2, V-B3 (PnP Solve), V-B4 (Joint Optical Flow) and V-B5 (3D Motion Refinement). (b) uses log-scale for the timing axes.
Fig. 10 presents the runtime breakdown for each module in the front-end. The feature tracking module, which includes feature extraction, matching, and geometric verification, requires less than $50\text{\,}\mathrm{ms}$. The motion estimation module, responsible for estimating the camera pose and all object motions per frame, averages $100\text{\,}\mathrm{ms}$ to produce all estimates. However, the majority of processing time is spent on the object motion refinement component, which takes approximately $\sim 250\text{\,}\mathrm{ms}$ per object to solve the nonlinear optimization problem. The runtime of this component depends heavily on the number of points involved in the optimization, leading to significant variation in processing times.
In our experiments, the time required for full-batch optimization ranges from $80\text{\,}\mathrm{s}$ to $700\text{\,}\mathrm{s}$, depending on the size of the constructed factor graph. While the full-batch optimization may not meet real-time requirements, it is important to note that practical applications typically rely on much smaller sliding window optimizations in the back-end. This approach significantly reduces computational demands and improves efficiency for online operation. In our experiments, using a window size of $20$, the average sliding-window optimization is $\sim 16\text{\,}\mathrm{s}$. Future work will focus on enhancing the efficiency of the back-end, including techniques such as conditional variable elimination [carlone2014eliminating], to better support real-time applications.
VII Downstream Tasks
@@
The proposed Dynamic SLAM system accurately estimates the motions of dynamic objects, as validated in Sec. VI. By eliminating the need for prior knowledge of motion models or object categories, our framework enables a range of downstream applications. This section explores how accurate motion estimation, produced by our pipeline, can facilitate dynamic object reconstruction and trajectory prediction, both of which are crucial for navigation systems.
VII-A Dynamic Object Reconstruction